A block of metal has a width of 3.2 cm

cm, a length of 17.1 cm

cm, and height of 5.8 cm

cm . Its mass is 1.3 kg

kg . Calculate the density of the metal.


1
Expert's answer
2021-09-02T02:02:08-0400

3.2 × 17.1 × 5.8= 317.376cm3

1.3 × 1000 = 1300

= 1300/317.376 = 4.096ml/cm3